OpenStudy (anonymous):

the area of a circle is found with the formula A=(pi)r^2, where pi represents 3.14. If a tile artists is laying down tiles in a circular mosaic pattern and the area of the circle 706.5 square feet, what is the radius of the circle in feet?

OpenStudy (anonymous):

706.5=3.14r^2 706.5/3.14=r^2 \[\sqrt{706.5\div3.14}\] =r^2

OpenStudy (anonymous):

=r not =r^2

OpenStudy (anonymous):

it says the choices are A. 11 B. 13 C. 15 D. 17

OpenStudy (anonymous):

its 15